OpenStudy (anonymous):

Please help me, I will give medal:) The value of y varies directly with x^2(x squared),and y=150 when x=5.What is the value of y when x=3?

OpenStudy (welshfella):

as this is direct variation we can write y = k x^2 where k is a constant

OpenStudy (anonymous):

ok, so would I square 3?

OpenStudy (welshfella):

given that y = 150 when x = 5:- 150 = k * 5^2 can you work out what k is ?

OpenStudy (anonymous):

ummmm, no :(

OpenStudy (welshfella):

leave the squaring of 3 until later the first part of the solution is to find the value of k

OpenStudy (welshfella):

5^2 = 25 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so that would be K?

OpenStudy (welshfella):

ok so we have 25k = 150 and to find k we divide both sides by 25 so k = 150 /25 = ?

OpenStudy (anonymous):

6

OpenStudy (welshfella):

right now we have the values of k k is called the constant of variation now we have established the relationship between y and x it is y = 6x ( replacing k by 6)

OpenStudy (welshfella):

now we can answer the quest ' What is y when x = 3

OpenStudy (welshfella):

oops sorry thats y = 6x^2 not 6x

OpenStudy (anonymous):

y=9?

OpenStudy (welshfella):

so when x = 3 y = 6 * (3)^2

OpenStudy (anonymous):

y=54

OpenStudy (welshfella):

right

OpenStudy (anonymous):

yay!thank you so much for helping me :D

OpenStudy (welshfella):

yw